# All Models Are Miscalibrated, But Some Less So: Comparing Calibration with Conditional Mean Operators

### v1

@@ Abstract
When working in a high-risk setting, having well calibrated probabilistic predictive models is a crucial requirement. However, estimators for calibration error are not always able to correctly distinguish which model is better calibrated. We propose the conditional kernel calibration error (CKCE) which is based on the Hilbert-Schmidt norm of the difference between conditional mean operators. By working directly with the definition of strong calibration as the distance between conditional distributions, which we represent by their embeddings in reproducing kernel Hilbert spaces, the CKCE is less sensitive to the marginal distribution of predictive models. This makes it more effective for relative comparisons than previously proposed calibration metrics. Our experiments, using both synthetic and real data, show that CKCE provides a more consistent ranking of models by their calibration error and is more robust against distribution shift.
@@ 1 Introduction
Probabilistic models quantify uncertainty by predicting a probability distribution over possible labels, rather than simply giving a point estimate. This allows models to express their confidence in a prediction. However, to ensure that the probability outputs can be trusted by users, such models must satisfy certain calibration (or reliability) properties. Roughly speaking, calibration requires that predicted probabilities match observed frequencies of labels. A classical example is in weather forecasting. A forecaster that assigns 70% chance of rain is well calibrated if, on average, it actually rains on 70% of days with such predictions. Although calibration was originally studied in a meteorological context (Brier,, 1950; DeGroot and Fienberg,, 1983), it has proven to be useful in a wide range of applications.
The viewpoint we take here is agnostic to whether the predictive model is Bayesian or frequentist and this should not in any way interfere with the essentially frequentist notion of calibration which we embrace. This is in the spirit of Little, (2006): “inferences under a particular model should be Bayesian, but model assessment can and should involve frequentist ideas.
@@
Work by Guo et al., (2017) cast a spotlight on the miscalibration of deep neural networks, which often give overconfident predictions. More recently, Phan et al., (2025) showed similarly poor calibration of state-of-the-art large language models. This gives reason to be concerned about the use of advanced AI models in high-risk settings. However, properly measuring calibration remains a challenge, with most authors and practitioners resorting to proxy metrics that do not fully capture the notion of strong calibration explained in Section 2.
@@
Reliability diagrams (DeGroot and Fienberg,, 1983), such as in Figure 1, offer a method of inspecting calibration for binary classifiers by comparing the model’s confidence in a positive result and its accuracy. For a well calibrated model, the confidence in a prediction should match the accuracy. While reliability diagrams provide a useful visual tool, they fail to quantify calibration error, and are difficult to adapt to multiclass and regression problems. A popular calibration metric inspired by reliability diagrams is the Expected Calibration Error (ECE) (Naeini et al.,, 2015) which measures the expected difference between the predictive confidence and observed frequencies of labels. But estimating ECE can be problematic as it is highly sensitive to the chosen binning scheme which leads to biased and inconsistent estimators (Nixon et al.,, 2019; Vaicenavicius et al.,, 2019; Widmann et al.,, 2019).
Figure 1: Reliability diagram for a binary classification problem comparing three probabilistic models along with the estimate of their conditional kernel calibration error (CKCE) where a lower value indicates better calibration.
@@
An alternative approach to calibration is via finding discrepancies between probability distributions, which can be represented by their kernel mean embeddings in a reproducing Hilbert space (RKHS). This idea was introduced by Widmann et al., (2021), who measured calibration by the difference in kernel embeddings of joint distributions of predictive distributions and labels, in order to construct a hypothesis test for calibration. While this is an effective tool for the purposes of testing, to which much work has been devoted (Widmann et al.,, 2021; Lee et al.,, 2023; Glaser et al.,, 2023; Chatterjee et al.,, 2024), in practice these tests may find that all sufficiently sophisticated models depart from perfect calibration. This motivates the idea of ranking a set of candidate models by their calibration error, similar to how models may be benchmarked by their accuracy or other metrics. However, as we demonstrate here, the calibration measure introduced by Widmann et al., (2021), as well as the classical ECE, both suffer from being impacted by the marginal distribution of the predictive distributions in a given model and can thus be ill-suited for comparing different candidate models.
@@
In this paper, we introduce the conditional kernel calibration error (CKCE). It evaluates calibration by determining the discrepancy between conditional distributions which we represent by their RKHS conditional mean operator. This can be viewed as complementary to the work of Widmann et al., (2021). It might be challenging to use CKCE as a test statistic for a calibration test because of the difficult to estimate null distribution, but we are interested in a different problem: reliable relative calibration comparisons between models. We demonstrate through a range of experiments on both synthetic and real data that the CKCE is better suited for this task.
@@ 2.1 Probabilistic models and Calibration
We begin with a description of probabilistic classifiers, followed by an extension to a more general setting. Suppose we aim to predict a label $y\in\mathcal{Y}=\{1,\dots,m\}$ based on a feature $x\in\mathcal{X}$. Rather than making a point prediction, we fit a probabilistic predictive model $Q$ which estimates the true conditional distribution $P(Y|X=x)\in\Delta^{m-1}$, where $\Delta^{m-1}$ represents the $(m-1)$-dimensional probability simplex. Thus, the model induces a probabilistic classifier $f:\mathcal{X}\to\Delta^{m-1}$, which maps a feature $x$ to the predicted conditional distribution of $Y$ given $X=x$ under $Q$, that is, $f(x)=Q(Y|X=x)$.
@@
$\mathbb{P}(Y=\arg\max f(X)|\max f(X)=c)=c$ (1)
for all $c\in[0,1]$ (Guo et al.,, 2017; Vaicenavicius et al.,, 2019). However, (1) does not consider calibration of all classes, which may be crucial if the entire predicted probability vector is used for decision making or as an intermediate output for downstream tasks, particularly when certain classes carry varying levels of risk for those tasks. Therefore, our focus is on the notion of so called strong calibration (Widmann et al.,, 2019).
Definition 2.1.
@@
$\mathbb{P}(Y=i|f(X)=q)=q^{(i)}\qquad\text{for }i=1,\dots m$ (2)
@@
We note that calibration is a property of a model that is essentially orthogonal to that of accuracy. Indeed, a predictive model which for any input simply returns the marginal distribution of class labels, $Q(Y|X)=P(Y)$, is calibrated according to Definition 2.1, but has accuracy at a chance level. Conversely, a classifier may attain the optimal (Bayes) risk with respect to the 0/1 loss (i.e. the highest possible accuracy) but have arbitrarily bad calibration. An example would be a classifier that assigns all probability mass in $Q(Y|X)$ to the label $\hat{Y}=\arg\max_{i=1,\ldots,k}P(Y=i|X)$.
Now consider a more general setting where labels $y$ belong to an arbitrary domain $\mathcal{Y}$, and denote by $\mathcal{P}$ the space of Borel probability measures over $\mathcal{Y}$. We can once again fit a probabilistic model $Q$, and define $Q_{X}=Q(Y|X)$ as a random variable taking a value in $\mathcal{P}$. Thus, the predictive model induces a joint distribution $P(Y,Q_{X})$ on the product space $\mathcal{Y}\times\mathcal{P}$, which allows us to generalise the notion of strong calibration.
Definition 2.2.
A probabilistic predictive model $Q_{X}=Q(Y|X)$ of the true conditional probability $P(Y|X)$ is strongly calibrated if
$P(Y|Q_{X})=Q_{X}$ (3)
almost surely $P(X)$.
The discrepancy between the left and right hand sides of (3) in Definition 2.2 can be understood as the calibration error of probabilistic model $Q$. Thus, evaluating calibration is ultimately a problem of measuring the discrepancy between two probability distributions. However, while $Q_{X}$ is readily available from the model output, $P(Y|Q_{X})$ can be highly complex, making it difficult to estimate. This motivates the use of nonparametric methods which allow to represent probability distributions and distances between them using only samples from the distributions.
@@ 2.2 Kernel Embeddings of Distributions
We provide a short review of some preliminaries on kernel embeddings of probability distributions in a reproducing kernel Hilbert space (RKHS) (Muandet et al.,, 2017). Consider a random variable $X$ with domain $\mathcal{X}$ and distribution $P(X)$. Let $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be any positive definite function with associated RKHS $\mathcal{H}_{k}$, which is a Hilbert space of functions $f:\mathcal{X}\to\mathbb{R}$ satisfying
  1. 1.
    For all $x\in\mathcal{X}$, $k(x,\cdot)\in\mathcal{H}_{k}$, and
  2. 2.
    For all $x\in\mathcal{X}$ and $f\in\mathcal{H}_{k}$, $\langle f,k(x,\cdot)\rangle_{\mathcal{H}_{k}}=f(x)$,
where 2 is known as the reproducing property. We can think of $k$ as the inner product between feature embeddings, that is, $k(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle_{\mathcal{H}_{k}}$ where $\phi(x)=k(x,\cdot)$ is the canonical feature map.
Assuming $k$ is bounded, the kernel mean embedding of $P(X)$ is given by
$\mu_{X}=\mathbb{E}_{X}k(X,\cdot)=\mathbb{E}_{X}\phi(X)\in\mathcal{H}_{k},$
and has the property $\langle\mu_{X},f\rangle_{\mathcal{H}_{k}}=\mathbb{E}_{X}f(X)$ for all $f\in\mathcal{H}_{k}$. Given another random variable $Y$ whose distribution has kernel embedding $\mu_{Y}$, we can compute the distance between $P(X)$ and $P(Y)$ using the maximum mean discrepancy (MMD) given by
$\text{MMD}_{k}^{2}(P(X),P(Y))$ $=\|\mu_{X}-\mu_{Y}\|_{\mathcal{H}_{k}}^{2}$
$=\mathbb{E}_{X,X^{\prime}}k(X,X^{\prime})+\mathbb{E}_{Y,Y^{\prime}}k(Y,Y^{\prime})$
$\qquad-2\mathbb{E}_{X,Y}k(X,Y),$
which can be easily estimated using samples from $P(X)$ and $P(Y)$ Gretton et al., (2012). For so called characteristic kernels, the mean embeddings are injective and the MMD equals zero if and only if the distributions $P(X)$ and $P(Y)$ are equal (Fukumizu et al.,, 2007). Many commonly used kernels such as the Gaussian and Laplace kernels are characteristic. A closely related notion is a universal kernel, which when defined on a compact metric space $\mathcal{X}$, has an associated RKHS $\mathcal{H}$ that is dense in the space of continuous functions on $\mathcal{X}$ (Sriperumbudur et al.,, 2011).
We can similarly define the kernel embedding of a joint distribution $P(X,Y)$ by choosing a kernel on the product domain $\mathcal{X}\times\mathcal{Y}$. Given kernels $k$ and $\ell$ on $\mathcal{X}$ and $\mathcal{Y}$ respectively, with associated RKHS $\mathcal{H}_{k}$ and $\mathcal{H}_{\ell}$, and defining the tensor product kernel $(k\otimes\ell)((x,y),(x^{\prime},y^{\prime}))=k(x,x^{\prime})\ell(y,y^{\prime})$, a special case for the joint mean embedding is given by
$\mu_{XY}=\mathbb{E}_{XY}[\phi(X)\otimes\psi(Y)]\in\mathcal{H}_{k\otimes\ell}$
where $\phi(x)=k(x,\cdot)$, $\psi(y)=\ell(y,\cdot)$ are the canonical feature maps (Fukumizu et al.,, 2004). By isometry between $\mathcal{H}_{k\otimes\ell}$ and $\mathcal{H}_{k}\otimes\mathcal{H}_{\ell}$, the joint mean embedding $\mu_{XY}$ can be identified with the uncentered cross-covariance operator $\mathcal{C}_{XY}:\mathcal{H}_{\ell}\to\mathcal{H}_{k}$, which has the property $\mathbb{E}[f(X)g(Y)]=\langle f,\mathcal{C}_{XY}g\rangle_{\mathcal{H}_{k}}$.
@@ 2.3 Conditional Mean Operators
Conditional distributions $P(Y|X=x)$ can likewise be embedded in an RKHS by the conditional mean embedding (Song et al.,, 2009)
$\mu_{Y|X=x}=\mathbb{E}_{Y|X=x}[\psi(Y)|X=x]\in\mathcal{H}_{\ell}.$
Note that the conditional mean embedding defines a family of points in the RKHS indexed by the values $x$. This motivates us to consider the conditional mean operator (CMO) $\mathcal{C}_{Y|X}:\mathcal{H}_{k}\to\mathcal{H}_{\ell}$ which satisfies $\mu_{Y|X=x}=\mathcal{C}_{Y|X}k(x,\cdot)$. Assuming that for all $g\in\mathcal{H}_{\ell}$, we have $\mathbb{E}_{Y}[g(Y)|X=\cdot]\in\mathcal{H}_{k}$, the CMO is a well defined Hilbert-Schmidt operator. Much research has been devoted to put conditional mean embeddings and operators on a rigorous footing, including the measure theoretic view taken by Park and Muandet, (2020). Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ sampled from $P(X,Y)$, an empirical estimator of the CMO is
@@
where indicates the adjoint of an operator, $\Psi_{Y}=[\psi(y_{1}),\dots,\psi(y_{n})]$, $\Phi_{X}=[\phi(x_{1}),\dots,\phi(x_{n})]$, $K_{XX}=\Phi_{X}^{*}\Phi_{X}$ is the Gram matrix with entries $[K_{XX}]_{ij}=k(x_{i},x_{j})$, $I_{n}$ is the $n\times n$ identity matrix and $\lambda$ is a regularisation parameter (Song et al.,, 2009). Assuming that $\mathcal{C}_{YX}\mathcal{C}_{XX}^{-3/2}$ is Hilbert-Schmidt, then $\|\widehat{\mathcal{C}}_{Y|X}-\mathcal{C}_{Y|X}\|_{HS}=O_{p}(\lambda^{1/2}+\lambda^{-3/2}n^{-1/2})$. Therefore, if the regularisation term satisfies $\lambda\to 0$ and $n\lambda^{3}\to\infty$, then $\widehat{\mathcal{C}}_{Y|X}$ converges in probability to the true CMO (Song et al.,, 2010, Theorem 1). A detailed refined analysis of convergence of conditional mean embeddings is studied in Li et al., (2022). Additionally, (Singh et al.,, 2023, Supplementary Material, Section 8) provides a minimax optimal rate for the convergence of CMO in Hilbert-Schmidt norm, which can depend on kernel $k$ and the smoothness of underlying conditional distributions. These results can provide a recipe for selecting the schedule of the regularisation parameter $\lambda$. An alternative method is to view the CMO as the solution to a regression problem, allowing $\lambda$ to be selected via cross-validation (Grünewälder et al.,, 2012). For simplicity, we choose the regularisation parameter to have schedule $\lambda_{n}=n^{-1/4}$.
Consider two random variables $Y$ and $Z$ with common domain $\mathcal{Y}$ and a conditioning random variable $X$. Assuming both CMOs are well defined, if $\mathcal{C}_{Y|X}=\mathcal{C}_{Z|X}$, then the conditional distributions $P(Y|X)$ and $P(Z|X)$ are equal in the sense that $P(Y|X=x)=P(Z|X=x)$ for $P(X)$-almost all $x\in\mathcal{X}$ (Ren et al.,, 2016, Theorem 3). Hence, the distance between conditional probability distributions is measured by the conditional maximum mean discrepancy (CMMD) (Ren et al.,, 2016) which is defined as
$\text{CMMD}_{k,\ell}^{2}(P(Y|X),P(Z|X))=\|\mathcal{C}_{Y|X}-\mathcal{C}_{Z|X}\|_{HS}^{2},$
where $k$ and $\ell$ are kernels on $\mathcal{X}$ and $\mathcal{Y}$ respectively. In practice, empirical CMO estimators from equation (4) can be used to estimate CMMD.

### v2

@@ Abstract
When working in a high-risk setting, having well calibrated probabilistic predictive models is a crucial requirement. However, estimators for calibration error are not always able to correctly distinguish which model is better calibrated. We propose the conditional kernel calibration error (CKCE) which is based on the Hilbert-Schmidt norm of the difference between conditional mean operators. By considering calibration error as the distance between conditional distributions, which we represent by their embeddings in reproducing kernel Hilbert space, the CKCE is less sensitive to the marginal distribution of predictive model outputs. This makes it more effective for relative comparisons than previously proposed calibration metrics. Our experiments, using both synthetic and real data, show that CKCE provides a more reliable measure of a model’s calibration error and is more robust against distribution shift.
@@ 1 Introduction
Probabilistic models quantify uncertainty by predicting a probability distribution over possible labels, rather than simply giving a point estimate. This allows models to express their confidence in a prediction. However, to ensure that the probability outputs can be trusted by users, such models must satisfy certain calibration (or reliability) properties. Roughly speaking, calibration requires that predicted probabilities match observed frequencies of labels. A classical example is in weather forecasting. A forecaster that assigns 70% chance of rain is well calibrated if, on average, it actually rains on 70% of days with such predictions. Although calibration was originally studied in a meteorological context [Brier_1950, DeGroot_1983], it has proven to be useful in a wide range of applications.
@@
Work by [Guo_2017] cast a spotlight on the miscalibration of deep neural networks, which often give overconfident predictions. More recently, [phan2025hle] showed similarly poor calibration of state-of-the-art large language models. This gives reason to be concerned about the use of advanced AI models in high-risk settings. However, properly measuring calibration remains a challenge, with most authors and practitioners resorting to proxy metrics that do not fully capture the notion of strong calibration explained in Section 2.
@@
Reliability diagrams [DeGroot_1983], such as in Figure 1, offer a method of inspecting calibration by comparing the model’s confidence in the predicted label and its accuracy, which should be equal for well calibrated forecasters. While reliability diagrams provide a useful visual tool, they fail to quantify calibration error, and are difficult to adapt to strong calibration. A popular calibration metric inspired by reliability diagrams is the expected calibration error (ECE) [naeini_2015] which measures the expected difference between the predictive confidence and observed frequencies of labels. But this quantity is heavily influenced by the marginal distribution of model predictions, making direct model comparison difficult. Furthermore, estimating ECE can be problematic as it is highly sensitive to the chosen binning scheme which leads to biased and inconsistent estimators [Nixon_2019, vaicenavicius_2019, Widmann_2019].
Calibration Error
ECE CKCE (ours)
ResNet 0.316 0.061
ViT 0.389 0.044
Figure 1: Reliability diagram and distribution of prediction confidence for two models trained on ImageNet. A lower CKCE value indicates a closer match between model confidence and accuracy, whereas ECE is heavily affected by the marginal distribution of predictions.
@@
An alternative approach to calibration is via finding discrepancies between probability distributions, which can be represented by their kernel mean embeddings in a reproducing kernel Hilbert space (RKHS). This idea was used by [Widmann_2021] to measure calibration via the difference in embeddings of joint distributions. While this is an effective tool for the purposes of a calibration hypothesis test, to which much work has been devoted [Widmann_2021, Lee_2023, Glaser_2023, Chatterjee_2024], the proposed estimator is once again sensitive to the marginal distribution of predictive confidence. Therefore, just like the classical ECE, the calibration error estimator proposed by [Widmann_2019] can be ill-suited for comparing different candidate models, particularly in situations with distribution shift.
@@
In this paper, we introduce the conditional kernel calibration error (CKCE). It evaluates calibration by determining the discrepancy between conditional distributions which we represent by their RKHS conditional mean operator. This can be viewed as complementary to the work of [Widmann_2021]. It might be challenging to use CKCE as a test statistic for a calibration test because of the difficulty in estimating the null distribution, but we are interested in a different problem: reliable relative calibration comparisons between models. We demonstrate through a range of experiments on both synthetic and real data that the CKCE is a more robust metric for this task.
@@ 2.1 Probabilistic Models and Calibration
Consider a classification problem in which we aim to predict a label $y\in\mathcal{Y}=\{1,\dots,m\}$ based on a feature $x\in\mathcal{X}$. Rather than making a point prediction, we fit a probabilistic classifier $f:\mathcal{X}\to\Delta^{m-1}$ which estimates the conditional distribution $P(Y|X=x)\in\Delta^{m-1}$, where $\Delta^{m-1}$ represents the (m1)(m-1)-dimensional probability simplex.
@@
$P(Y=\arg\max f(X)|\max f(X)=c)=c$ (1)
for all $c\in[0,1]$ [Guo_2017, vaicenavicius_2019]. However, (1) does not consider calibration of all classes, which may be crucial if the entire predicted probability vector is used for decision making or as an intermediate output for downstream tasks, particularly when certain classes carry varying levels of risk for those tasks. Therefore, our focus is on the notion of so called strong calibration [Widmann_2019].
Definition 1
@@
$P(Y=i|f(X)=q)=q^{(i)}\qquad\text{for }i=1,\dots m$ (2)
@@
We note that calibration is a property of a model that is essentially orthogonal to that of accuracy. Indeed, a predictive model which for any input simply returns the marginal distribution of class labels, $f(x)=P(Y)$, is calibrated according to Definition 1, but has accuracy at a chance level. Conversely, a classifier may attain the optimal (Bayes) risk with respect to the 0/1 loss (the highest possible accuracy) but have arbitrarily bad calibration. An example would be a classifier that assigns all probability mass to label $\hat{Y}=\arg\max_{i=1,\ldots,m}P(Y=i|X)$.
If we let $Q_{X}=f(X)$, then the discrepancy between the probability distributions $P(Y|Q_{X})$ and $Q_{X}$ (corresponding to the left and right hand sides of (2) in Definition 1) can be understood as the calibration error of the probabilistic model. However, while $Q_{X}$ is readily available from the model output, $P(Y|Q_{X})$ can be highly complex, making it difficult to estimate. This motivates the use of nonparametric methods which allow to represent probability distributions and distances between them using only samples from the distributions.
@@ 2.2 Kernel Embeddings of Distributions
We provide a short review of some preliminaries on kernel embeddings of probability distributions in a reproducing kernel Hilbert space (RKHS) [muandetKernelMeanEmbedding2017a]. Consider a random variable $X$ with domain $\mathcal{X}$ and distribution $P(X)$. Let $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be any positive definite function with associated RKHS $\mathcal{H}_{k}$ and canonical feature map $\phi(x)=k(x,\cdot)$.
Assuming $k$ is bounded, the kernel mean embedding of $P(X)$ is given by $\mu_{X}=\mathbb{E}_{X}k(X,\cdot)\in\mathcal{H}_{k}$ and has the property $\langle\mu_{X},f\rangle_{\mathcal{H}_{k}}=\mathbb{E}_{X}f(X)$ for all $f\in\mathcal{H}_{k}$. Given another random variable $Y$ with embedding $\mu_{Y}$, we can compute the distance between $P(X)$ and $P(Y)$ using the maximum mean discrepancy (MMD)
$\text{MMD}_{k}^{2}(P(X),P(Y))$ $=\|\mu_{X}-\mu_{Y}\|_{\mathcal{H}_{k}}^{2}$ (3)
$=\mathbb{E}_{X,X^{\prime}}k(X,X^{\prime})+\mathbb{E}_{Y,Y^{\prime}}k(Y,Y^{\prime})-2\mathbb{E}_{X,Y}k(X,Y)$
The MMD can be easily estimated using samples from $P(X)$ and $P(Y)$ [Gretton_2012]. For so called characteristic kernels, the mean embeddings are injective and the MMD equals zero if and only if the distributions $P(X)$ and $P(Y)$ are equal [Fukumizu_2007]. Many commonly used kernels such as the Gaussian and Laplace kernels are characteristic. A closely related notion is a universal kernel, which when defined on a compact metric space $\mathcal{X}$, has an associated RKHS $\mathcal{H}$ that is dense in the space of continuous functions on $\mathcal{X}$ [sriperumbudurUniversalityCharacteristicKernels2011].
We can similarly define the kernel embedding of a joint distribution $P(X,Y)$ by choosing a kernel on the product domain $\mathcal{X}\times\mathcal{Y}$. Given kernels $k$ and $\ell$ on $\mathcal{X}$ and $\mathcal{Y}$ respectively, with associated RKHS $\mathcal{H}_{k}$ and $\mathcal{H}_{\ell}$, and defining the tensor product kernel $(k\otimes\ell)((x,y),(x^{\prime},y^{\prime}))=k(x,x^{\prime})\ell(y,y^{\prime})$, a special case for the joint mean embedding is $\mu_{XY}=\mathbb{E}_{XY}[\phi(X)\otimes\psi(Y)]\in\mathcal{H}_{k\otimes\ell}$ where $\phi(x)=k(x,\cdot)$, $\psi(y)=\ell(y,\cdot)$ are the canonical feature maps [Fukumizu_2004]. By isometry between $\mathcal{H}_{k\otimes\ell}$ and $\mathcal{H}_{k}\otimes\mathcal{H}_{\ell}$, the joint mean embedding $\mu_{XY}$ can be identified with the uncentered cross-covariance operator $\mathcal{C}_{XY}:\mathcal{H}_{\ell}\to\mathcal{H}_{k}$, which has the property $\mathbb{E}[f(X)g(Y)]=\langle f,\mathcal{C}_{XY}g\rangle_{\mathcal{H}_{k}}$.
@@ 2.3 Conditional Mean Operators
Conditional distributions $P(Y|X=x)$ can likewise be embedded in an RKHS by the conditional mean embedding $\mu_{Y|X=x}=\mathbb{E}_{Y|X=x}[\psi(Y)|X=x]\in\mathcal{H}_{\ell}$ [Song_2009]. Note that the conditional mean embedding defines a family of points in the RKHS indexed by the values $x$. This motivates us to consider the conditional mean operator (CMO) $\mathcal{C}_{Y|X}:\mathcal{H}_{k}\to\mathcal{H}_{\ell}$ which satisfies $\mu_{Y|X=x}=\mathcal{C}_{Y|X}k(x,\cdot)$. Assuming that for all $g\in\mathcal{H}_{\ell}$, we have $\mathbb{E}_{Y}[g(Y)|X=\cdot]\in\mathcal{H}_{k}$, the CMO is a well defined Hilbert-Schmidt operator. Much research has been devoted to put conditional mean embeddings and operators on a rigorous footing, including the measure theoretic view taken by [Park_2020]. Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ sampled from $P(X,Y)$, an empirical estimator of the CMO is
@@
where indicates the adjoint of an operator, $\Psi_{Y}=[\psi(y_{1}),\dots,\psi(y_{n})]$, $\Phi_{X}=[\phi(x_{1}),\dots,\phi(x_{n})]$, $K_{XX}=\Phi_{X}^{*}\Phi_{X}$ is the Gram matrix with entries $K_{XX}]_{ij}=k(x_{i},x_{j})$, $I_{n}$ is the $n\times n$ identity matrix and $\lambda$ is a regularisation parameter [Song_2009]. Assuming that $\mathcal{C}_{YX}\mathcal{C}_{XX}^{-3/2}$ is Hilbert-Schmidt, then $\|\widehat{\mathcal{C}}_{Y|X}-\mathcal{C}_{Y|X}\|_{HS}=O_{p}(\lambda^{1/2}+\lambda^{-3/2}n^{-1/2})$. Therefore, if the regularisation term satisfies $\lambda\to 0$ and $n\lambda^{3}\to\infty$, then $\widehat{\mathcal{C}}_{Y|X}$ converges in probability to the true CMO [Song_2010, Theorem 1]. A detailed refined analysis of convergence of conditional mean embeddings is studied in [Li_2022]. Additionally, [singh_2023, Supplementary Material, Section 8] provides a minimax optimal rate for the convergence of CMO in Hilbert-Schmidt norm, which can depend on kernel $k$ and the smoothness of underlying conditional distributions. These results can provide a recipe for selecting the schedule of the regularisation parameter $\lambda$. An alternative method is to view the CMO as the solution to a regression problem, allowing $\lambda$ to be selected via cross-validation [Grunewalder_2012]. For simplicity, we choose the regularisation parameter to have schedule $\lambda_{n}=n^{-1/4}$.
Consider two random variables $Y$ and $Z$ with common domain $\mathcal{Y}$ and a conditioning random variable $X$. Assuming both CMOs are well defined, if $\mathcal{C}_{Y|X}=\mathcal{C}_{Z|X}$, then the conditional distributions $P(Y|X)$ and $P(Z|X)$ are equal in the sense that $P(Y|X=x)=P(Z|X=x)$ for $P(X)$-almost all $x\in\mathcal{X}$ [Ren_2016, Theorem 3]. Hence, the distance between conditional probability distributions is measured by the conditional maximum mean discrepancy (CMMD) [Ren_2016] which is defined as
$\text{CMMD}_{k,\ell}^{2}(P(Y|X),P(Z|X))=\|\mathcal{C}_{Y|X}-\mathcal{C}_{Z|X}\|_{HS}^{2},$ (5)
where $k$ and $\ell$ are kernels on $\mathcal{X}$ and $\mathcal{Y}$ respectively, and $\|\cdot\|_{HS}$ represents the Hilbert-Schmidt norm of an operator. In practice, empirical CMO estimators from equation (4) can be used to estimate the CMMD.